OpenStudy (anonymous):

Which could be the next step in simplifying this expression? 13 + 19 + 7 A. 20 + 19 (use the commutative property) B. 32 + 20 (use the distributive property) C. 13 + 36 (use the associative property) D. 36 + 13 (use the commutative property)

OpenStudy (anonymous):

@Utterly_Confuzzled

OpenStudy (utterly_confuzzled):

I'd say a based upon the properties and the rules followed

OpenStudy (anonymous):

?????

OpenStudy (anonymous):

@Utterly_Confuzzled i dont understand

OpenStudy (utterly_confuzzled):

Okay, so in this problem, distributive property is completely irrelevant, so you can mark out B. Associative property is (a+b)+c = a+(b+c), which doesn't happen in the answers. That narrows it down to A or D. D ends up being wrong because 7+19 is 26, not 36. That leaves A as your answer

OpenStudy (anonymous):

so a is my answer

OpenStudy (utterly_confuzzled):

Correct

OpenStudy (anonymous):

thx for helping me!

OpenStudy (utterly_confuzzled):

No problem at all!!

OpenStudy (anonymous):

Translate the word phrase into a math expression. three less than the product of six and a number

OpenStudy (anonymous):

can you help me in this one too?

OpenStudy (anonymous):

A. 3<6n B. 6/n - 3 C. 6n - 3 D. 3 - 6n

OpenStudy (utterly_confuzzled):

Yes It would be 6x-3 because that is 3 less than 6 times a number

OpenStudy (utterly_confuzzled):

So C

OpenStudy (anonymous):

plz hurry i think i have to go soon

OpenStudy (anonymous):

so C

OpenStudy (utterly_confuzzled):

Right

OpenStudy (anonymous):

The length of a rectangular field is 4 m longer than its width. If the perimeter is 56 m, what is the width of the field? A. 8 m B. 12 m C. 14 m D. 16 m

OpenStudy (anonymous):

@Utterly_Confuzzled

OpenStudy (utterly_confuzzled):

56=(2x)+(2x +8) 48=4x x=12 12+4=16 D

OpenStudy (anonymous):

so its D

OpenStudy (utterly_confuzzled):

Yes

OpenStudy (anonymous):

Felipe rode 5600 meters on his bike. How many kilometers did he ride? A. 5.6 km B. 560 km C. 56,000 km D. 5,600,000 km

OpenStudy (utterly_confuzzled):

1000meters = 1 km 5600/1000=?

OpenStudy (anonymous):

5.6

OpenStudy (utterly_confuzzled):

Right!!

OpenStudy (anonymous):

Which inequality correctly orders –28,12, –19, and 2? A. –28 < 12 < –19 < 2 B. –28 < –19 < 2 < 12 C. –19 < –28 < 2 < 12 D. 2 < –19 < –28 < 12

OpenStudy (anonymous):

@Utterly_Confuzzled

OpenStudy (utterly_confuzzled):

B It basically reads (-28) is less and (-19) is less than 2 is less than 12

OpenStudy (anonymous):

Solve. –258 – (–307 ) + (–361) A. –926 B. –312 C. –204 D. 410

OpenStudy (anonymous):

Solve. –258 – (–307 ) + (–361) A. –926 B. –312 C. –204 D. 410

OpenStudy (anonymous):

@Utterly_Confuzzled

OpenStudy (utterly_confuzzled):

-258 - (-307) + (-361) -258 + 307 - 361 Can you solve it from there?

OpenStudy (anonymous):

is it B?

OpenStudy (utterly_confuzzled):

Yups!

OpenStudy (anonymous):

thx a lot! :)

OpenStudy (utterly_confuzzled):

No problem! Anytime!